# Resolving Task Objective Conflicts in Unified Multimodal Understanding and Generation via Task-Aware Mixture-of-Experts

### v1

@@
Abstract
Unified multimodal large language models (MLLMs) based on end-to-end autoregressive (AR) transformers effectively integrate both understanding and generation tasks within a single framework. However, intrinsic Task Objective Conflicts between high-level semantic abstraction in understanding and fine-grained detail preservation in generation pose significant challenges, often leading to suboptimal trade-offs and task interference. Existing solutions, such as decoupling shared visual encoders, fall short of fundamentally resolving these conflicts due to inherent AR architecture. In this paper, we propose a novel approach that decouples internal components of AR to resolve task objective conflicts. Specifically, we design UTAMoE, a Unified Task-Aware Mixture-of-Experts (MoE) framework that decouples internal AR modules via a Task-Aware MoE Layer to create task-specific optimization subpaths. To enhance task differentiation while maintaining overall coordination, we introduce a novel Two-Stage Training Strategy. Extensive experiments on multimodal benchmarks demonstrate that UTAMoE mitigates task objective conflicts, achieving state-of-the-art performance across various tasks. Visualizations and ablation studies further validate the effectiveness of our approach.
Resolving Task Objective Conflicts in Unified Multimodal Understanding and Generation via Task-Aware Mixture-of-Experts
Jiaxing Zhang Sichuan University zjxing9972@gmail.com Xinyi Zeng Sichuan University perperstudy@gmail.com Hao Tang††thanks: Corresponding author. Peking University haotang@pku.edu.cn
1 Introduction
Recently, the emergence of end-to-end autoregressive (AR) Transformer architectures Wu et al. (2024c); Xie et al. (2024) have emerged as an effective solution for unified modeling of multimodal understanding and generation, enabling genuine integration of both tasks within a single framework. However, a fundamental challenge arises from the intrinsic divergence in task objectives: understanding tasks aim to distill high-level semantic representations, whereas generation tasks require preserving and integrating fine-grained details. This task objective conflict often leads to interference during training, making it difficult for models to effectively balance both tasks.
Recent efforts, such as Janus Wu et al. (2024a); Ma et al. (2024); Chen et al. (2025), attempt to mitigate this conflict by decoupling the shared visual encoder, enabling specific visual representations and reducing feature-level interference. However, such strategies do not fundamentally resolve the divergence in task objectives. As illustrated in Figure 1 a), current models struggle to achieve an optimal trade-off between understanding and generation. This is reflected in issues like semantic degradation in comprehension (e.g., answering "What is on the grass?" with a generic label "animal"), or missing fine-grained details in generation (e.g., omitting the "puppy" in "A boy sits on a chair with a puppy beside him").
The encoder-level decoupling proposed by Janus motivates our further exploration of whether internal modules within the autoregressive architecture can be decoupled to better balance conflicting objectives across tasks. In contrast to encoder-level decoupling, internal decoupling requires a careful trade-off between maintaining architectural unification and enabling task-specific specialization. Excessive decoupling may undermine shared contextual understanding and parameter efficiency, while insufficient decoupling may fail to alleviate cross-task interference. Therefore, a mechanism is needed to enable task-specific while maintaining unified training and computational efficiency.
To this end, we identify Mixture of Experts (MoE) Dai et al. (2024); Liu et al. (2024a) as a promising strategy for structural decoupling. By dynamically activating different experts to construct task-specific subpaths, MoE allows disentangling feature spaces and reducing cross-task interference—thereby enabling efficient and specialized processing in unified multimodal architectures.
In this paper, we propose UTAMoE, a Unified Task-Aware Mixture-of-Experts framework designed to address Task Objective Conflicts in multimodal understanding and generation. The key idea is to construct task-oriented optimization subpaths within AR using a Task-Aware MoE Layer. Additionally, we propose a Two-Stage Training Strategy to further enhance task-specific differentiation while maintaining overall performance balance. As shown in Figure 1 b), UTAMoE yields more desirable outcomes in both tasks—accurate recognition (e.g., "puppy") in understanding and improved detail preservation in generation.
We conduct extensive experiments on multiple multimodal benchmarks and observe that UTAMoE achieves substantial improvements in semantic recognition accuracy for understanding tasks, along with enhanced detail preservation for generation tasks. Both quantitative, qualitative, and ablation studies validate the effectiveness of decoupling the internal AR architecture in mitigating task objective conflict. We hope that UTAMoE offer novel perspectives for unified multimodal tasks and offer valuable insights into effectively addressing task objective conflicts identified in this work.
Our main contributions are summarized below:
We propose a novel approach to decouple internal components of autoregressive (AR) architectures, effectively mitigating Task Objective Conflicts.
@@ 1 Introduction
We present UTAMoE, a Task-Aware MoE framework that enables task-specific optimization within unified multimodal models.
We introduce a Two-Stage Training Strategy that helps UTAMoE achieve state-of-the-art performance across multiple benchmarks.
2 Related Work
Mixture of Experts (MoE).
@@ 2 Related Work
Recent studies Dai et al. (2024); Liu et al. (2024a); Cai et al. (2024) have shown MoE’s effectiveness when integrated with language models. For instance, Mixtral-MoE 8x7B Jiang et al. (2024) excels in scaling model capacity and improving task performance. DeepSeek introduced Top-K routing and shared experts, enhancing MoE’s task allocation efficiency. Variants like Chen et al. Chen et al. (2023b) enable multitask learning for simple tasks through embedding analysis, while Li et al. Li et al. (2025) assign different experts for different modalities, proving expert specialization’s feasibility. These advancements strongly support MoE’s integration into unified multimodal frameworks.
Multimodal Understanding.
@@
Multimodal large language models (MLLMs), such as LLaVA Liu et al. (2023), MiniGPT-4 Zhu et al. (2023), and InstructBLIP Panagopoulou et al. (2023), are strong in processing multimodal information. They project features from pretrained modality-specific encoders (e.g., CLIP Radford et al. (2021)) into LLMs to enable cross-modal reasoning. However, these models focus mainly on visual perception and lack native generative ability. Recent studies Ge et al. (2023); Jin et al. (2023); Sun et al. (2023) explore extending MLLMs to other domains, which presents an area worth further exploration.
Visual Generation.
Visual generation is a rapidly evolving field where diffusion-based methods have excelled in text-to-image and text-to-video tasks. These methods operate in a continuous latent space obtained via a VAE encoder, predicting Gaussian noise added to the latent representations. Inspired by natural language processing, autoregressive (AR) models use transformer architectures to predict discrete visual tokens Ramesh et al. (2021); Esser et al. (2021). These models discretize visual data and apply GPT-style mechanisms for generation. Additionally, masked prediction models adopt BERT-like strategies to improve generation efficiency by predicting masked input portions. Such approaches have been extended to more complex multimodal tasks, including video generation Yu et al. (2023a).
@@
Unified multimodal understanding and generation models are regarded as highly effective for enabling seamless reasoning and content generation across diverse modalities Team (2024); Zhou et al. (2024). However, existing approaches, whether based on autoregressive (AR) architectures with shared encoders Wu et al. (2024c); Team (2024) or decoupled ones Wu et al. (2024a); Ma et al. (2024); Chen et al. (2025), often suffer from task objective conflicts when jointly addressing both understanding and generation tasks. To address this conflict, our UTAMoE leverages the MoE mechanism to establish task-specific optimization pathways within AR architecture, thereby effectively addressing task objective conflict.
3 The Proposed Method
Recently, Unified multimodal large language models based autoregressive (AR) Transformers Team (2024); Wu et al. (2024c) address understanding and generation tasks in a single framework. However, the conflicts arising from the fundamental divergence between the objectives of understanding and generation tasks reveal the inherent limitations of the autoregressive (AR) architecture. Janus decouples the shared visual encoder to produce task-specific visual representations, effectively mitigating interference at the visual feature level across tasks. However, this approach does not fully resolve the task objective conflict. To address this issue, we propose a novel approach centered on decoupling the internal components of the autoregressive (AR) architecture. Specifically, we design a Task-Aware MoE Layer, as detailed in subsection 3.1, to replace the conventional feed-forward network (FFN) layer, thereby forming our unified architecture—UTAMoE. This module incorporates two key mechanisms: Hierarchical Expert Routing for task-specific expert selection and Hybrid Expert Collaboration for inter-task synergy.
Additionally, to fully leverage unique strengths of different expert modules and improve overall generalization of the framework, we propose a novel two-stage training strategy, detailed in subsection 3.2, to further enhance performance of UTAMoE.
3.1 Task-Aware MoE Layer
@@
To better accommodate task objective conflict of understanding and generation tasks, we introduce a modular design based on a Mixture-of-Experts (MoE) framework. Specifically, the expert are divided into two groups, each tailored to understanding and generation tasks, respectively. To enhance task-specific specialization, we propose a Hierarchical Expert Routing mechanism, which comprises a task-aware router and a conventional dynamic assignment router. This mechanism enables more accurate routing of task-relevant tokens to the appropriate expert. Furthermore, to ensure coherence and synergy across task objectives, we incorporate a shared expert independent of the task-specific groups. As a core part of the Hybrid Expert Collaboration mechanism, the shared expert facilitates cross-task information exchange and helps maintain a balanced representation between understanding and generation tasks.
3.1.1 Hierarchical Expert Routing
@@
Task-Aware Router.
Given an input token representation , we first classify it into one of the two task-specific expert groups via a softmax-based classifier:
@@
Dynamic-Assignment Router.
Once the expert group is determined, we further compute the relevance between the token and each expert in the group, and select the top- experts based on the scores:
@@
3.1.2 Hybrid Expert Collaboration
@@
where is a learnable hyperparameter that balances the contributions.
@@
where is the ground-truth for the task group, is the loss for task , and , are hyperparameters controlling the loss weights.Specifically, corresponds to the understanding task, while controls the generation task.
3.2 Training Strategy
To fully leverage the unique capabilities of different expert modules and enhance the generalization ability of the overall framework, we propose a Two-stage Training strategy, as illustrated in Figure 3, to construct UTAMoE. In the following paragraphs, we detail the objectives and provide a comprehensive explanation of each training stage.
Stage 1: Training Task-Specific Experts.
This stage focuses on acquiring task-specific expert through dedicated training on individual tasks. The objective is to obtain expert weights that perform well on their respective tasks. Specifically, we adopt the generative entropy loss as the primary training criterion. During this process, the self-attention layers are frozen, and only the feed-forward networks (FFNs) are updated, resulting in well-specialized task-specific experts.
Stage 2: LoRA-Based Fine-Tuning of UTAMoE.
In this stage, we first replace the conventional FFN layers with the Task-Aware MoE layer. Then, we integrate the previously obtained expert weights and perform joint fine-tuning of UTAMoE using a mixture of understanding and generation tasks. Specifically, the Task-Aware MoE layer utilizes a task-aware router to dynamically select the appropriate task group, within which a Top-K strategy is applied to choose task-specific experts. We then perform end-to-end fine-tuning using the LoRA approach to optimize the entire model while maintaining high parameter efficiency.
Through LoRA fine-tuning, our UTAMoE efficiently adapts to diverse tasks while maintaining high parameter efficiency, thereby enhancing the model’s generalization capability and enabling effective training in multi-task settings.
4 Experiments
In this section, we introduce comprehensive experiments to evaluate our method.Firstly, we outline our experimental setup, including model architecture and training configurations. Next, we evaluate our model on various understanding and generation benchmarks. We also provide qualitative results to further illustrate the strengths of our approach. Finally, we analyze the expert utilization for both task-specific expert and shared expert.
4.1 Experimental Setups
@@
We adopt JanusPro as the base model, whose language component supports a maximum sequence length of 4096. For multimodal understanding, we use SigLIP-Large-Patch16-384 Zhai et al. (2023) as the image encoder. For visual generation, the encoder utilizes a codebook of size 16,384 and downsamples inputs by a factor of 16. Both understanding and generation adapters are implemented as two-layer MLPs. Additionally, we integrate the Any-Resolution technique based Liu et al. (2024d) to support inputs at varying resolutions.
@@
Since our training data are subsets of those used in Wu et al. (2024a) and we use JanusPro as base model, the vision tower weights are kept frozen throughout all training stages, as they have already been sufficiently pretrained and generalize well across similar data distributions. For training task-specific experts, we utilize ShareGPT4V Chen et al. (2024b) dataset for understanding tasks and MidJourney Vivym (2023) dataset for generation tasks. During the LoRA-based fine-tuning stage, a mixed dataset combining both sources is employed. Regarding the MoE configuration, each task group is assigned two experts, and a Top-K routing strategy is applied within the group. Additionally, inspired by DeepSeekMoE Dai et al. (2024), one shared expert is included across all groups to encourage generalization. All experiments are conducted on a single NVIDIA V100 GPU (32GB). We set , , and , with the remaining hyperparameters detailed in Table 1.
| Hyperparameter | Stage 1 | Stage 2 |
| Batch size | 2 | 2 |
| Learning rate | ||
| Optimizer | AdamW (, ) | |
| LR Scheduler | Cosine | |
| Weight decay | 0.0 | 0.0 |
| Training steps | 200 | 400 |

### v2

# Resolving Task Objective Conflicts in Unified Model via Task-Aware Mixture-of-Experts

@@
Abstract.
Recently, multimodal understanding (MMU) and text-to-image generation (T2I) have been integrated into a single autoregressive (AR) architecture, achieving initial unification. However, existing works focus on representation-level studies and overlook potential conflicts in AR architectures’ internal information flow during training different tasks. Motivated by this gap, we identify a deeper issue, Task Objective Conflict (TOC), arising from AR architectures’ internal information flow, which causes negative transfer and catastrophic forgetting when training MMU and T2I jointly. To address this issue, we proposed UniDecouple, which decouples internal modules for different tasks to construct task-specific optimization subpaths. To implement UniDecouple, we employ a Task-Aware Mixture-of-Experts (TA-MoE), comprising Hierarchical Expert Routing and Hybrid Expert Collaboration, trained in two stages: first to build task-specific experts, then jointly fine-tuned to balance specialization and overall coordination. Extensive experiments on both understanding and generation benchmarks demonstrate that UniDecouple preserves strong understanding ability while achieving generation quality comparable to state-of-the-art methods, offering a new perspective for unified multimodal modeling.
Key words and phrases:
Unified Model, Autoregressive, Mixture-of-Experts1. Introduction
Human cognition, where understanding and generation intricately interleave, provides the foundation for reasoning and thought Islam and Moushi (2025); Wu et al. (2024a); Tong et al. (2024). Inspired by this mechanism, researchers have sought unified models that integrate multimodal understanding and generation within a single framework. Yet progress remains divided: autoregressive (AR) based models Zhu et al. (2024b); Liu et al. (2024d); Zhu et al. (2023); Bai et al. (2023) drive multimodal understanding (MMU), while diffusion approaches Podell et al. (2023); Wang et al. (2024c); Rombach et al. (2022) dominate text-to-image generation (T2I). To bridge these divergent paradigms, recent works (e.g., DALL·E Reddy et al. (2021)) employ AR based models for T2I, resulting in a preliminarily unified paradigm and prompting further investigations Team (2024); Chern et al. (2024); Dong et al. (2023); Wang et al. (2024b); Wu et al. (2024c) into extending AR frameworks to support both MMU and T2I. Yet, a well-balanced trade-off between enhancing generation quality and retaining understanding ability remains challenging for AR based models.
A widely recognized challenge for AR based unified models arises from representation level Wu et al. (2024a); Ma et al. (2024); Chen et al. (2025). MMU relies on semantic abstraction to capture high-level information, whereas T2I demands fine-grained detail preservation to faithfully render the image. The conflict in representation requirements has been widely recognized, and to reconcile it, early approaches first adopted a single pixel or semantic encoder to learn shared representations Wu et al. (2024c); Lin et al. (2025). Nevertheless, relying on a single encoder often biases the extracted features toward one task and cannot adequately satisfy these distinct objectives. Subsequently, JanusPro Chen et al. (2025) decouples the encoders, leveraging both pixel and semantic encoders to derive task-specific representations, effectively resolving the representation conflict.
Nevertheless, prior work on unifying MMU and T2I within AR based architectures focuses on surface-level representation conflicts, overlooks the deeper issues inherent in the architecture itself. Consequently, we focus on the internal information flow and propose that conflicts also emerge within the architecture Sener and Koltun (2018); Yu et al. (2020). Subsequently, we conduct theoretical analyses and empirical experiments to validate this hypothesis, demonstrating that these conflicts, termed Task Objective Conflict (TOC), indeed exist. Specifically, training MMU and T2I simultaneously using AR based approaches can lead to negative transfer, where learning on one task (e.g., understanding) inadvertently degrades performance on another (e.g., generation), and catastrophic forgetting, where previously acquired task-specific knowledge is partially lost.
To mitigate TOC, we propose UniDecouple. As illustrated in Fig. 1(c), UniDecouple uniquely integrates with the AR paradigm by leveraging a Mixture of Experts (MoE), whose intrinsic task-specific routing mechanism provides dedicated subpaths for each task. Nevertheless, the conventional MoE architecture remains inadequate for well resolving these conflicts, as it lacks explicit mechanisms to orchestrate task-specific experts and reconcile the divergent objectives of MMU and T2I. To this end, we design Task-Aware MoE (TA-MoE), which incorporates key components: Hierarchical Expert Routing, enabling more precise task-specific pathway selection, and Hybrid Expert Collaboration, facilitating the integration of complementary knowledge. We also introduce a two-stage training strategy: training experts for task-specific skills and jointly fine-tuned to balance specialization and overall coordination.
1
We empirically validate the effectiveness of UniDecouple through a series of experiments. In Sec. 4.2, we first validate our hypothesis by analyzing the behavior of the loss function and examining how TOC manifests during training. In Secs. 4.3, we then conduct both quantitative and qualitative evaluations on mainstream understanding and generation benchmarks, demonstrating that UniDecouple consistently achieves strong performance across tasks. Finally, in Sec. 5, ablation studies confirm the contributions of our proposed TA-MoE architecture and the two-stage training strategy. Overall, these results demonstrate the robustness and effectiveness of UniDecouple. Our main contributions are summarized below:
We validate the existence of Task Objective Conflict (TOC) in widely adopted autoregressive (AR) frameworks from both theoretical and empirical perspectives.
@@ Abstract.
We propose UniDecouple, a Task-Aware MoE (TA-MoE) framework that enables task-specific optimization within unified multimodal models.
Our comprehensive experiments show that UniDecouple achieves generation quality on par with current state-of-the-art methods while effectively preserving the model’s understanding ability.
2. Related Work
Mixture of Experts (MoE)
@@ 2. Related Work
Recent studies Dai et al. (2024); Liu et al. (2024b); Cai et al. (2024) have demonstrated the effectiveness of MoE when integrated with large language models. For instance, Mixtral-MoE 8x7B Jiang et al. (2024) excels in scaling model capacity and improving task performance. DeepSeek introduced Top-K routing and shared experts, enhancing MoE’s task allocation efficiency. Variants such as Chen et al. Chen et al. (2023a) enable multitask learning for simple tasks through embedding analysis, while Li et al. Li et al. (2025) assign different experts for different modalities, proving the feasibility of expert specialization.
Multimodal Understanding (MMU)
@@
Multimodal large language models (MLLMs), such as Flamingo Alayrac et al. (2022), LLaVA Liu et al. (2023), MiniGPT-4 Zhu et al. (2023), BLIP-2 Li et al. (2023b), InstructBLIP Panagopoulou et al. (2023), Qwen2-VL Wang et al. (2024a), and Intern-VL Chen et al. (2024c), demonstrate strong capabilities in processing multimodal information. They typically align features from pretrained modality-specific encoders (e.g., CLIP Radford et al. (2021)) using methods such as Q-Former and connectors, and then map them into the LLM to enable cross-modal reasoning. Recent studies Ge et al. (2023); Jin et al. (2023); Sun et al. (2023) have attempted to extend MLLMs beyond perception to generation and other domains, opening new research directions.
Text-to-Image Generation (T2I)
Generation has undergone a series of methodological advancements. Initially, diffusion-based methods generated high-quality samples by iteratively denoising Gaussian perturbations in continuous latent spaces Ho et al. (2020); Nichol and Dhariwal (2021). The traditional denoising process relied on U-Net, it has gradually evolved to Transformer-based prediction method Peebles and Xie (2023). Diffusion models excel in generation fidelity and stability, but their multi-step iterative process results in slow inference. To improve efficiency, flow matching Lipman et al. (2022) generative models were introduced, learning continuous probability flows to more directly approximate the data distribution while retaining the advantages of diffusion. Subsequently, autoregressive (AR) models began to tackle visual generation tasks Reddy et al. (2021), discretizing image into token sequences and employing Transformer to predict each token sequentially. These advancements lay a solid foundation for the development of subsequent unified multimodal models.
@@
Unified multimodal understanding and generation models, based on AR, are regarded as highly effective for enabling seamless reasoning and content generation across diverse modalities Team (2024); Zhou et al. (2024). However, existing approaches, whether based on autoregressive (AR) architectures with shared encoders Wu et al. (2024c); Team (2024) or decoupled ones Wu et al. (2024a); Ma et al. (2024); Chen et al. (2025), often suffer from task objective conflicts (TOC) when jointly addressing both understanding and generation tasks. These conflicts can lead to negative transfer and also slow down convergence during joint training. To address this conflict, our UniDecouple leverages the MoE mechanism to establish task-specific optimization pathways within AR architecture, thereby effectively addressing TOC.
2
3. UniDecouple
In this section, we first provide a theoretical analysis of the Task Objective Conflicts (TOC) in Section 3.1, which highlight the inherent limitations of unified autoregressive (AR) frameworks when addressing both MMU and T2I tasks. Building on this foundation, we introduce the fundamental component of our proposed Unidecouple architecture: Task-Aware Mixture of Experts (TA-MoE) Layer, which decouples the internal structure to enable task-specific expert selection and effective cross-task collaboration (details presented in Section 3.2). Furthermore, to fully exploit the complementary strengths of different expert modules and enhance the overall generalization capability of the framework, we propose a two-stage training strategy in Section 3.3.
3.1. Task Obejective Conflict
We formalize the TOC within the context of AR modeling, highlighting how sequential dependency and attention masking amplify gradient conflict between tasks.
Consider an AR model parameterized by , trained on a fixed dataset to perform two tasks: MMU and T2I. The model predicts a token sequence under an auto-regressive factorization:
where is modeled by causal attention. The two tasks differ in objective scope: MMU abstracts semantics from observed tokens, while T2I reconstructs unobserved tokens over longer horizons, with task losses defined as:
MMU mainly updates early tokens () for representation abstraction, while T2I gradients flow from later tokens () for detailed reconstruction. Under sequential training (understanding first, generation second), parameter updates are:
with denoting the learning rate. Because both losses propagate through overlapping attention pathways, their gradients are not independent but correlated through shared causal attention weights:
where is the Jacobian of hidden states w.r.t. , and denotes token-level prediction error. When the semantic abstraction (MMU) and detailed reconstruction (T2I) objectives induce opposite attention gradients on overlapping tokens, the inner product
becomes negative, revealing a gradient-level conflict specific to AR architectures, where the resulting degradation in understanding loss after generation updates can be approximated as:
Because under negative inner products, the model exhibits negative transfer and catastrophic forgetting.
3.2. Task-Aware MoE Layer
@@
To better accommodate task objective conflict of understanding and generation tasks, we introduce a modular design based on a MoE framework. Specifically, the expert are divided into two groups, each tailored to understanding and generation tasks, respectively. To enhance task-specific specialization, we propose a Hierarchical Expert Routing mechanism, which comprises a task-aware router and a conventional dynamic assignment router. This mechanism enables more accurate routing of task-relevant tokens to the appropriate expert. Furthermore, to ensure coherence and synergy across task objectives, we incorporate a shared expert independent of the task-specific groups. As a core part of the Hybrid Expert Collaboration mechanism, the shared expert facilitates cross-task information exchange and helps maintain a balanced representation between MMU and T2I tasks.
3.2.1. Hierarchical Expert Routing
@@
Task-Aware Router. Given an input token representation , we first classify it into one of the two task-specific expert groups via a softmax-based classifier:
@@
Dynamic-Assignment Router. Once the expert group is determined, we further compute the relevance between the token and each expert in the group, and select the top- experts based on the scores:
@@
3.2.2. Hybrid Expert Collaboration
@@
where is a learnable hyperparameter to balance contributions.
@@
where is the ground truth for the task group, is the loss for task , and , are hyperparameters controlling the loss weights. Specifically, corresponds to the understanding task, while controls the generation task.
3.3. Training Strategy
To fully leverage the unique capabilities of different experts and enhance the generalization ability of the overall framework, we propose a Two-stage Training strategy, as illustrated in Figure 3. In the following paragraphs, we detail objectives and provide a comprehensive explanation of each training stage.
3
Stage 1: Training Task-Specific Experts. This stage concentrates on deriving task-specific experts via specialized training for each individual task. The aim is to obtain expert weights that excel in their corresponding tasks. Mathematically, we employ the generative entropy loss as the main training objective, which can be formulated as:
where denotes the true probability distribution of the i-th sample over C classes, and is the predicted probability from the task-specific expert. During this training phase, the self-attention layers, characterized by their parameter set , are frozen (i.e., ), and only the feed-forward networks (FFNs) with parameter are updated. This leads to the formation of highly specialized task-specific experts, as expressed by the parameter update rule for the FFN:
where is the learning rate and represents the training iteration.
Stage 2: LoRA-Based Fine-Tuning of UniDecouple. This stage, we first replace the traditional FFN layers with the Task-Aware MoE (TAMoE) layer. The TAMoE layer’s output can be described as:
Here, are the weights assigned by the task-aware router to select the top- experts from either the understanding group (Und. Group) or the generation group (Gen. Group), and is the output of the -th expert in group for input . Then, we incorporate the expert weights obtained in Stage 1 and conduct joint fine-tuning of UniDecouple using a combination of understanding and generation tasks.
Specially, we perform end-to-end LoRA fine-tuning with a rank of , enabling UniDecouple to efficiently adapt to diverse tasks with improved generalization through highly parameter-efficient optimization.
4. Experiments
In this section, we present comprehensive experiments for our proposed method. First, we describe the experimental setup, including the model architecture and training configurations. Next, we verify the existence of task objective conflicts (TOC) through metrics such as loss and evaluate the model on various understanding and generation benchmarks. We then provide qualitative results to illustrate the strengths of our approach. Finally, we analyze the utilization of both task-specific and shared experts.
4.1. Experimental Setups
@@
We adopt JanusPro as the base model, whose language component supports a maximum sequence length of 4096. For multimodal understanding, we use SigLIP-Large-Patch16-384 Zhai et al. (2023) as the image encoder. For image generation, the encoder utilizes a codebook of size 16,384 and downsamples inputs by a factor of 16. Both understanding and generation adapters are implemented as two-layer MLPs.
@@
For the training setup, we keep the vision tower frozen in all stages since it is sufficiently pretrained by Chen et al. (2025). Task-specific experts are trained on ShareGPT4V Chen et al. (2024b) for understanding and MidJourney Vivym (2023) for generation, while the LoRA fine-tuning stage uses a mixed dataset of both. In the MoE design, each task group is assigned two experts with Top- routing, and one shared expert is added to enhance generalization. The hyperparameters are as follows: in Stage 1, batch size of 32, learning rate , and 200 training steps; in Stage 2, batch size of 32, learning rate , and 400 training steps. Both stages use the AdamW optimizer (), cosine learning rate scheduler, and weight decay of 0.0. The main loss coefficients are set to , , and .
4.2. Existence of Conflict
First, we design two variants based on Chen et al. (2025): (1) a single-task model, where each task is trained independently, and (2) a multi-task model, where task-specific losses are jointly optimized. Using POPE and GenEval to cover understanding and generation tasks, results show that the multi-task model suffers performance drops on certain tasks and fails to maintain balanced performance, demonstrating inherent conflicts between task objectives.
| Task | POPE | GenEval | POPE | Gen.Eval | |
| (a) | Und. | 86.9 | – | – | – |
| (b) | Gen. | – | 0.74 | – | – |
| (c) | Both | 86.2 | 0.73 | -0.7 | -0.01 |
Second, to further verify TOC, we analyze the dynamics of the loss functions during multi-task training. Specifically, we track the loss values of multimodal understanding and image generation tasks over training iterations to identify whether decreases in one task’s loss correspond to increases in another’s. Our understanding task uses cross-entropy loss, while the generation task employs mean squared error (MSE) loss. If there is significant mutual suppression in loss curves, it indicates the existence of optimization conflicts between tasks.
From Figure 4, we observe a certain degree of antagonism between the two loss functions, which further supports the existence of task conflicts. However, these experiments do not provide a thorough verification of the conflict, and addressing this at a deeper level remains a goal for our future work.
Loss
4.3. Quantitative Evaluation
Multimodal Understanding (MMU). To assess MMU, we evaluate our model on widely recognized understanding benchmarks, which include POPE Li et al. (2023a), MME Fu et al. (2024), MMB Liu et al. (2024a), SEED Li et al. (2024a), GQA Hudson and Manning (2019), MMMU Yue et al. (2024) and MM-Vet Yu et al. (2023).
Text-to-Image Generation (T2I). For evaluating T2I capabilities, we use GenEval Ghosh et al. (2023) and DPG-Bench Hu et al. (2024).
Comparison Methods. As shown in Table 2 and Table 4, we categorize the models into three groups based on their capabilities. (i) Understanding Only: This group includes models designed solely for MMU, such as the LLaVA series (e.g., LLaVA Liu et al. (2023), LLaVA-v1.5 Liu et al. (2024c)), the MobileVLM series (e.g., MobileVLM Chu et al. (2023), MobileVLM-V2 Chu et al. (2024)), InstructBLIP Dai et al. (2023), Qwen-VL-Chat Bai et al. (2023), and Emu3-Chat Wang et al. (2024c). (ii) Generation Only: This group contains models that focus exclusively on image generation, including LlamaGen Sun et al. (2024), LDM Rombach et al. (2022), SDv1.5 Rombach et al. (2022), PixArt- Chen et al. (2023b), SDv2.1 Rombach et al. (2022), DALL·E 2 Ramesh et al. (2022), Emu3-Gen Wang et al. (2024c), SDXL Podell et al. (2023), DALL·E 3 Betker et al. (2023), and SD3-Medium Esser et al. (2024). (iii) Understanding and Generation: This group comprises models capable of both MMU and T2I, such as DreamLLM Dong et al. (2023), LaViT Jin et al. (2023), MetaMorph Tong et al. (2024), NEXT-GPT Wu et al. (2024b), Show-o-256 Xie et al. (2024), SDHo-512 Xie et al. (2024), D-DiT Li et al. (2024b), Gemini-Nano-1 Team et al. (2023), ILLUME Wang et al. (2024b), TokenFlow-XL Qu et al. (2024), LWM Liu et al. (2024e), VILA-U Wu et al. (2024c), Chameleon Team (2024), and the Janus-series models Wu et al. (2024a); Chen et al. (2025); Ma et al. (2024).